# Unlocking Mental Health: Exploring College Students’ Well-being through Smartphone Behaviors

### v1

@@ III-B Unlocking Behavior and Mental Well-being Correlation
Overall Correlations. To answer RQ2, we conduct correlation analysis as described in Section II-D and the results are shown in Table I. Notably, all p-values for the correlation coefficients are well below 0.01, indicating strong associations between phone usage patterns and mental health. This finding highlights the potential of our novel unlocking features as effective predictors of mental health outcomes. Column 1 reveals an intriguing pattern: longer phone usage durations correlate with poorer mental health outcomes, whereas higher unlock frequencies are associated with better mental health. These findings underscore the importance of considering both frequency and duration when assessing the impact of smartphone use on mental health, and call for further studies to identify the root causes of these interesting phenomena.
@@ III-C Insights from Multinomial Logistic Regression
Gender Differences in Predictive Patterns. Among males, longer unlock durations are associated with a higher likelihood of being in the “Severe” mental health category, whereas for females, longer durations are more likely to be linked to “Normal” state rather than “Severe”. Notably, male students exhibit greater sensitivity to unlocking behaviors with a strong probability linking longer unlock duration to “Severe” mental status (coefficient>0.5). This reinforces our earlier observations regarding gender differences, highlighting the critical role of phone usage in male students’ mental health, as it appears to significantly increase the risk of severe stress and anxiety. These preliminary findings on the gender differences suggest that the mechanisms linking phone usage to mental health may vary significantly across groups. Consequently, predictive models for mental health based on phone behaviors should be tailored for specific groups—or even individuals—to capture these variations effectively.
@@ V Conclusion and Future Directions
This study presents the first large-scale investigation of college students’ smartphone unlocking behaviors and their association with mental well-being, leveraging over 210,000 data points from a four-year longitudinal dataset. Our findings highlight significant gender and contextual differences, emphasizing the need for fine-grained analyses in mental health research. We demonstrate that unlocking behaviors alone offer strong predictive power, enabling lightweight models for practical adoption. Future work should refine predictive models beyond gender differences, incorporate temporal dynamics such as high-stress periods, and develop personalized, context-aware interventions, including location-based recommendations, to promote healthier phone usage and digital well-being.

### v2

@@ III-B Unlocking Behavior and Mental Well-being Correlation
Overall Correlations. To answer RQ2, we conduct correlation analysis as described in Section II-D and the results are shown in Table I. Notably, all p-values for the correlation coefficients are well below 0.01, indicating associations exist between phone usage patterns and mental health. This finding highlights the potential of our novel unlocking features as effective predictors of mental health outcomes. Column 1 reveals an intriguing pattern: longer phone usage durations correlate with poorer mental health outcomes, whereas higher unlock frequencies are associated with better mental health. These findings underscore the importance of considering both frequency and duration when assessing the impact of smartphone use on mental health, and call for further studies to identify the root causes of these interesting phenomena.
@@ III-C Insights from Multinomial Logistic Regression
Gender Differences in Predictive Patterns. Among males, longer unlock durations are associated with a higher likelihood of being in the “Severe” mental health category, whereas for females, longer durations are more likely to be linked to “Normal” state rather than “Severe”. Notably, male students exhibit greater sensitivity to unlocking behaviors with a higher probability linking longer unlock duration to “Severe” mental status (coefficient>0.5). This reinforces our earlier observations regarding gender differences, highlighting the critical role of phone usage in male students’ mental health, as it appears to significantly increase the risk of severe stress and anxiety. These preliminary findings on the gender differences suggest that the mechanisms linking phone usage to mental health may vary significantly across groups. Consequently, predictive models for mental health based on phone behaviors should be tailored for specific groups—or even individuals—to capture these variations effectively.
@@ V Conclusion and Future Directions
This study presents the first large-scale investigation of college students’ smartphone unlocking behaviors and their association with mental well-being, leveraging over 210,000 data points from a four-year longitudinal dataset. Our findings highlight significant gender and contextual differences, emphasizing the need for fine-grained analyses in mental health research. We demonstrate that unlocking behaviors alone offer predictive power, enabling lightweight models for practical adoption. Future work should refine predictive models beyond gender differences, incorporate temporal dynamics such as high-stress periods, and develop personalized, context-aware interventions, including location-based recommendations, to promote healthier phone usage and digital well-being.